### v2

# Sorting-based FPGA Sliding Window Aggregation Engine without off-chip Memories

Abstract.
Aggregation queries are a series of computationally-demanding analytics operations on grouped and time series data. They include tasks such as summation or finding the median among the items of a group sharing a group ID, and within a specified number of the last observed tuples for sliding window aggregation (SWAG). They have a wide range of applications including in database analytics, operating systems, bank security and medical sensors. Existing challenges include the hardware complexity that comes with efficiently handling per-group states using hash-based approaches. This paper presents a pipelined and adaptable approach for calculating a wide range of aggregation queries with high throughput. It is then adapted for SWAG to achieve up to 476x speedup over the CPU of the same platform. It outperforms the state-of-the-art such as by being able to process 7.14x more tuples per second, and support 4x the window sizes with a fraction of the resources and no DRAM.
FPGA, reconfigurable, group-by-aggregate, engine, prefix scan, sliding window, SWAG, key-value pairs
1. Introduction
Aggregation is the task of combining raw data to produce more useful and summarised knowledge. It appears in many disciplines, especially relating to distributed and parallel programming models, such as MapReduce (Tangwongsan et al., 2018), where multiple workers/nodes produce a partial result that then needs aggregating. The worker notion is also popular in FPGA designs (Owaida et al., 2017). Equivalently, an established software design pattern is the fork-join model, which is a prominent concurrency pattern with processes or threads calculating partial results (De Wael et al., 2014).
In data science, there are numerous operations that can contribute to the usefulness of these computational models. One way is with the column and key-value-store paradigms, which are also increasingly adopted in high-performance analytics databases, as opposed to row-store. This is because of the potential to stream whole columns, which increases data localities and linearises the memory access pattern for certain analytics operations, also benefiting FPGA acceleration.
One common aggregation application is the group-by aggregate operator for relational databases. It provides a summary of rows belonging to (having) the same group ID. An example SQL (structured query language) snippet is provided in algorithm 1, where a sum operation is applied to all keys (table0.key2) of the same group (table0.key1). Other than summation, SQL aggregate functions include minimum, maximum, average and count. These are sometimes combined with more advanced options such as for calculating the distinct key count per group ID.
A similar category of queries is sliding window aggregation (SWAG) (Tangwongsan et al., 2018). A time domain is associated with the data, as with queries specifying a time range to form windows for aggregation. Due to the fact that the window formation is applied on-the-fly, near where the data is gathered, the stream is considered partly sorted with regard to accompanying timestamps or the implied order of the data. SWAG usually yields a random memory access pattern to calculate, as with fine-tuned CPU algorithms (Theodorakis et al., 2018). An example is when finding the median, as even with FPGAs, current designs require caching of larger states (Geethakumari and Sourdis, 2023). This complicates the design by increasing the on-chip and off-chip memory requirements for supporting/approaching the worst case (as with multiple inputs hashing to the same slot for hashsets).
1
2SELECT table0.key1, sum(table0.key2)
3FROM table0
4GROUP BY table0.key1
5ORDER BY table0.key1
Algorithm 1 SQL example of group-by-aggregate for summation.
This paper presents a novel design for a group-by-aggregate engine that integrates computation inside high-throughput and efficient structures such as the prefix scan. This engine is then adapted for performing SWAG queries, by adding additional functionality to an accompanying sorter. The proposed SWAG engine is highly memory-efficient to the point that it achieves equivalent functionality to highly complex competing designs without their main requirement for an off-chip DRAM (Ramakrishnan Geethakumari et al., 2019; Geethakumari and Sourdis, 2021a, b, 2023) or compression (Geethakumari and Sourdis, 2023). The internal state fits inside on-chip memories such as block-RAM. The main idea behind this efficiency is to replace the need for large indexed or hashed states by offloading the design complexity to small-scale sorting, as it eliminates any fragmentation that would be caused by the need to maintain per-group information.
The contributions of this paper are as follows:
  • An efficient adaptable group-by-aggregate engine for sorted streams that processes multiple inputs per cycle.
  • A state-of-the-art high-throughput DRAM-less sliding window aggregation (SWAG) engine based on the group-by-aggregate engine.
  • A series of modifications for a high-throughput sorter to be able to append group cardinality to each tuple for the purposes of supporting the SWAG median operator.
  • System-based and out-of-context evaluation of the proposed designs.
2. Background
An emerging need in data analytics is the use of programming models and related infrastructure to be able to benefit from the increasingly-wide memories and datapaths (Papaphilippou et al., 2023). A more traditional alternative is to have multiple workers and ports, as with FPGA-based database designs with multi-threading-equivalent arrangements. However, since the memory access pattern is highly-affected by this design choice, the hardware complexity is also affected, such as by co-designing cache or hash structures on the programmable logic.
FPGAs are more than capable of accelerating complex tasks such as aggregation while utilising wide datapaths. This can be achieved “naturally” by introducing state machines that perform fine-grain parallelisation, or with multiple-workers on buffered data. In order to simplify the design of such high-throughput operators, the hardware programmer can implement modular designs based on optimised building blocks.
Figure 1. Modular design of the distinct count query (right) with example data (Papaphilippou et al., 2020) in orange, based on the PRRA implementation (left) (Papaphilippou et al., 2023).
One recent example of a throughput-optimised building block is the reverse butterfly-based “parallel round-robin arbiter” (PRRA (Papaphilippou et al., 2023)), which removes sparsity from high-throughput data. It is essentially a round-robin load balancer, but it retains the order of the data to facilitate computation. Figure 1 illustrates how a group-by aggregation task (distinct count per group ID) can be built in a modular way using a PRRA. One of the aims of the presented work is further simplification of such pipelines for group-by aggregation, as we show that they have certain amounts of logic redundancy (as with the implementation discussion of section 4.5).
Figure 2. SWAG example for the minimum operator.
The main target of the presented work is sliding-window aggregation (SWAG). This family of queries operate on streaming data, and perform a filtering step that evaluates a statement relating to time to form windows for aggregation. The input mainly data consists of (group, key) tuples, also namely key-value pairs. The time is stated either by the natural order of their appearance in the input, or by appended timestamps (Oge et al., 2013). The scope of our work only regards the time being represented by the order of the tuples as they arrive. An example SWAG query is usually of the form “calculate the median of the last WS elements per group ID every WA tuples” (Geethakumari and Sourdis, 2023), where median can be replaced by any supported aggregate operator, WS is the window size, and WA is the distance between the sampled windows. Figure 2 shows an example for the minimum operator, where on every 2 tuples (WA) we find the minimum key per group within the last 4 tuples (WS).
3. Proposed solution
Figure 3. Five-step pipeline of the proposed group-by-aggregate engine.
In this section, we first propose a novel design for performing a common analytics set of queries (group-by-aggregate, section 3.1). This engine is then adapted to perform sliding window aggregation queries (SWAG, section 3.2). In order to support applicability to a wider range of query operators including median, an accompanying sorter is extended with group-cardinality functionality (section 3.3). The main theme of these designs is high-throughput, as they can accept $P$ elements per FPGA cycle, where $P$ is the degree of parallelism, aiming at scalability for saturating wide datapaths.
3.1. Group-by-aggregate engine
The proposed group-by-aggregate engine is a high-throughput pipeline for calculating a wide range of group-by-aggregate queries. It also works as a non-blocking building block, as it never asserts backpressure on the stream (i.e. the “ready” signal is high). Based on the query classification from (Tangwongsan et al., 2018), the supported types are Sum-like, Max-like and Sketch-like.
Figure 3 introduces the architecture of the proposed group-by-aggregate engine. A rolling prefix scan is used to calculate the permutation indices for the permutation network (reverse butterfly (Papaphilippou et al., 2023)) to achieve stream compaction. Here, the engine exploits the prefix-scan topology to calculate and append aggregation information to the data, before filtering redundant tuples (already aggregated). The engine expects a sorted stream consisting of ($group_{i},key_{i}$) tuples, arriving in batches of size $P$ per cycle. The input is assumed dense, as the engine either receives $P$ or 0 tuples in a cycle, according to other latencies at the system-level. In its output, the aggregation engine can create sparsity, hence the usefulness of the compaction functionality of the permutation network.
Steps (a) and (b)
The stage (a) in the pipeline is responsible for buffering up to one extra batch of size $P$, so that the entities (nodes) $t$ are able to (b) distinguish and mark each last tuple per $group$. Essentially, this marking is done by setting their $last$ bit high, representing the last tuple for each group. These last tuples are important, because they will be those that will carry the aggregate results, as they will be able to “see” all previous values of their $group$ before being extracted. Thus, these bits will also be used as valid-bits for the reverse butterfly network, to denote a valid tuple, also containing the summarised information after the end of the pipeline.
Step (c)
The entities named $n$ are distributed inside the adapted prefix scan to calculate the permutation index of the tuples to achieve the round-robin effect by the reverse butterfly network, as with the PRRA (Papaphilippou et al., 2023). At the same time they also operate on the $key$ fields, which later becomes the $result$ field containing the aggregation summary per group. For instance, if the query function is to find the sum of all the keys for each of the group IDs, the entities $n$ not only perform a cumulative sum for the reverse butterfly $index$, but also for the key sum per group, until the last entry of the same group is processed to carry the full sum of the group. An equality check on the group field with the previous entry ($group_{i}\stackrel{{\scriptstyle?}}{{=}}group_{i-1}$) is enough for knowing where not to propagate the previous result. Using sum as an example, the start of a new $group$ would leave the current $key$ intact instead of adding the result of the last tuple.
In order to also support the distinct count operator (as in figure 1), the entities $n$ also need to follow a simple distributed algorithm: if the propagated minimum from the input node of the larger key differs from the smaller key, then the distinct counts are coming from disjoint sets, hence $n$ adopts their sum. If the minimum is equal to the smaller key of the two, then $n$ adopts this sum of distinct counts, but subtracts 1 for taking care of the double counting of the common key.
Step (d)
The entities $n^{\prime}$ are special versions of the entities $n$ and their main functionality is to output the originally-marked tuples as valid to the permutation network, while replacing their key field with the result. The main difference with $n$ is that sometimes the partial result information cannot be contained in the $key$ field, and they are the entrypoint for taking additional signals into account for the final result.
For instance, if $function\_select$ indicates that the operation is to find the average values per group, the entities $n$ operate the same as in the sum operation. However, it is the entity $n^{\prime}$ that will divide the result by the corresponding group tuple count (cardinality). In such a case, the entities $n$ also include internal signals for counting the number of tuples in the group in a similar fashion to the sum operation, but adding 1 instead of the key. A related difference is with the count signals of $n^{\prime}$, since $n^{\prime}$ are responsible for the rolling effect, the propagated count of the bottom $n^{\prime}$ is able to count beyond $P$ elements, e.g. by having a 32-bit width. All $n^{\prime}$ elements use this value plus the $(log_{2}P+1)$-bit wide count from the prefix sum to provide the final 32-bit count in the result, when it is requested such as from a count query. According to the complexity of the supported operations, additional signals may be carried in the network, such as the minimum, that becomes of use when calculating the distinct count per group ID.
Step (e)
Finally, the reverse butterfly network permutes the batch of results to achieve the round-robin effect in parallel, and is inline with (Papaphilippou et al., 2023). The reverse butterfly network is used as an optimal-complexity switch, though other permutation networks can also be used. The $last$ flag is now renamed to a $valid$ bit, as it now indicates that a port has valid result.
Figure 4. Colour-coded numerical example: finding the average per group.
Figure 5. High-throughput sliding window aggregation. The main functionality is to select the min., median and max. values per group ID.
Figure 4 introduces a numerical example on how a single batch of $P$ tuples is processed through the pipeline for $P=4$. Initially, the 4 tuples have group IDs ranging from {a,b,c}, and they are sorted. The last tuple for each group is known for at least the first 3, plus the remaining one because of step $(a)$. In this example, it is easy to see that (a,2) and (b,4) are marked with T for true as the last elements with that group ID (the example assumes that the next batch also starts with c, hence (c,9) becomes (c,9,F) on the next pipeline stage. Then, the marked tuples pass through the prefix scan, which appends to them (incrementally): a cumulative sum on their values (to facilitate sum-related operations), a cumulative sum on their valid bits (to calculate the number of tuples per group), and a cumulative sum on their recently-marked last bits (that will be used as an index to the reverse butterfly network for the round-robin effect). The first two sums are in the scope of their corresponding group IDs, and it is achieved by simply comparing their group IDs before interacting with each other according to the prefix scan topology. The next step is the additional stage in the prefix scan, where the $n^{\prime}$ entities finalise the results and also update the counts according to the previous batch. Here, (a,2) and (b,4), being selected as the tuples to carry the aggregated information for their groups a and b respectively, they replace their keys with the average values per group (e.g. by dividing 12 by 2 for b). The permutation indices are not altered in this case, remaining at 0 and 1 respectively, assuming this was the first batch, or that the previous batch has outputted its last aggregate result on index $P-1$.
3.2. Sliding window aggregation engine
The design from section 3.1 is adapted to perform sliding-window aggregation (SWAG) queries. The main difference from the non-SWAG engine is that the input is bounded by windows. The context of the input has a fixed user-defined window size (WS), and a simple buffering mechanism provides these windows, while skipping by window advance (WA), as explained in the background (section 2). Additionally, instead of relying on an external sorter, a sorter with a special functionality needs to be part of the SWAG engine.
Step (a)
[…]
Step (b)
A sorter is needed right before the aggregation engine to sort the window based on the $group$ IDs, as well as the keys. Note that sorting by the keys as well is not always required, such as when the operator is commutative as with summation. Since 4K elements, for example, are considered moderately large as SWAG window sizes (Geethakumari and Sourdis, 2023), the sorting can entirely be done on-the-fly instead of indexing larger memories. FPGA-based high-throughput sorters are thoroughly studied in literature (Saitoh et al., 2018; Zuluaga et al., 2016; Qiao et al., 2022; Song et al., 2016).
Steps (c) and (d)
These steps remain similar as with the design of section 3.1, but with one of the three simultaneous cumulative sums (for keys, valid bits and last bits) decoupled from the others, in order to happen first. Hence, it needs two prefix scan-like structures. The preceding prefix scan is on the valid bits, so that the index of each tuple within the group is known first. The idea is to be able to do selection-based queries such as the median value, just by comparing this index with the group cardinality that is already-appended to all tuples by the modified sorter (section 3.3). In this work, we focus on the min./med./max. operator, which leaves the first, middle, and last tuples of each group (corresponding to the min./med./max.) as a valid aggregate result, as opposed to a single tuple for the aggregated summary of the group-by engine. If queries such as the median are not needed, the group-by engine of section 3.1 is sufficient to perform its simpler operators in a sliding-window context.
Figure 6. SWAG numerical example: finding the minimum, median and maximum value per group per sliding window.
Steps (e) and (f)
The last two steps reduce back to a PRRA architecture (Papaphilippou et al., 2023), as the remaining functionality only relates to the round-robin compaction. The final index calculation is done by the last prefix scan in the pipeline, while the first one is now responsible for all computation-related functionality including the other two cumulative sums. See the numerical example of figure 6, which has similar semantics as figure 4. It demonstrates the min./med./max. selection, as well as the expected positions of the results in the output after the abstracted steps (e) and (f) as a PRRA.
3.3. Sorter with group-cardinality calculation
Additional logic is needed when supporting non-incremental aggregation functions such as when finding the median. The calculation of the median value is not trivial without knowing the size of the data for which it is applied. This requires a modification inside the preceding sorter to be able to append the median-related information, which is the group cardinality alongside the data passed on to the aggregation engine. This functionality can conveniently be integrated into the sorter, as the sorter is provided with the entirety of the windows before flushing its sorted output to the aggregation engine.
The ability to involve median calculation within a sorter depends on the sorter architecture (Guo and Luk, 2001). Our proposed solution is based on a sorter that uses the linear sorter as a building block (Papaphilippou et al., 2020). The linear sorter is the hardware equivalent of insertion sort, as it is able to insert a new value into an already sorted sublist (Koch and Torresen, 2011). Each sorting cell, as depicted in figure 7 (center), has a comparator that compares its tuple with the incoming one. According to the result of the comparison and the one of its neighbour (“left smaller”), it either adopts the new value, propagates the left value, or remains as is.
Figure 7. Adding group-cardinality support to the linear sorter.
The modification that appends group cardinalities exploits linear sorter’s ability to have a sorted list on each cycle by comparing with every existing element. An additional comparison is done on the group subfield of the incoming tuple to determine if it is from the same group, and increments its own cardinality accordingly, as illustrated in figure 7. The incoming tuple then adopts the group cardinality upon insertion according to the neighbouring entity that holds the same group ID. It is guaranteed that a neighbour or its past value will have this information, since the group IDs appear clustered due to the sorting effect.
The baseline sorter of our design is able to provide high-throughput (multiple tuples per cycle) while supporting moderately large window sizes. The open-source sorter of (Papaphilippou et al., 2020) is specifically selected because it combines linear sorters with a merge tree (PMT (Song et al., 2016)) for high-throughput. This combination features special adaptations to be able to work efficiently in many-leaf (input list) merge mode using the same hardware.
Figure 8. Data movement inside the proposed SWAG engine, according to the window size (WS) and internal size ($k$) of the merge-capable sorter.
Figure 8 shows how the data moves throughout our proposed SWAG engine that uses this sorter. When the windows fit inside the sorter ($\textit{WS}\leq k$, where $k$ is the capacity of the sorter (i.e. the total cells of the linear sorters), the sorter works in sorting mode. In this mode, the linear sorters are flushed and then merged on-the-fly each time a window is read in its entirety. The output is then directly connected to the aggregation engine, which contains the sorted windows with each tuple having appended the new cardinality information. When $\textit{WS}>k$, the window is sorted in chunks of size $k$ and are stored temporarily in a BRAM buffer. The sorted chunks then reenter the sorter, but this time the sorter works in merge mode, and its linear sorters are able to perform merging by only keeping the minimum head of each sublist alongside the corresponding BRAM indices. During the merge phase, the modified sorter cells work as before, but the cardinality of incoming data is inherited from the stored values, instead of being equal to 1. Internally, this second pass makes our SWAG engine to be able to support window sizes (WS) of up to $k^{2}$.
Figure 9. Adding group-cardinality support to the sorter.
An additional modification is needed to propagate the correct group cardinalities, as summarised in figure 9. The parallel merge tree (PMT) needs to sum the cardinalities calculated by the linear sorters. Each FLiMS merger (Papaphilippou et al., 2022) of the PMT uses a network of compare-and-swap (CAS) units, which are sorters of 2-elements. The modification here is that each CAS unit uses the cond1 and cond2 conditions to determine how to update two incoming cardinalities every time they share the same group ID. The tuples carry two extra bits inside the mergers, one for denoting if the updating of their cardinalities has already happened, and one for their original source (lists A or B according to figure 9). Cond1 (to sum both cardinalities) is true whenever the cardinality update has not already happened to any of the 2 inputs, and the source is different. Cond2 (to adopt the other cardinality) is true whenever the cardinality update has happened to the other tuple, and the source is either different or the same. A similar arrangement is done for the selector stage units, as they are based on the same comparison, but they also remember any merged cardinalities from the elements of the last comparison. The latter is needed to support the case when a set of same group IDs is only coming from one of the lists A or B, such as when the ones from the other list have already been extracted.
The idea of a proof for this algorithm relates to the fact that the tuples sharing the same group ID (most significant bits) will appear consecutively in the output of the FLiMS merger, since its output is always sorted. As the order of those “duplicates” in regard to their group ID will also be correct in regard to their other bits, this means the required pairwise comparisons take place to establish this order, which are enough to broadcast such information, as with spanning trees (Touzene and Day, 2004).
4. Evaluation
This section evaluates the implementations of the group-by aggregate and SWAG engines. The studied pipelines accept a data stream consisting of $(group,key)$ pairs. The engines are highly parameterisable, but the evaluation mostly focuses on the following configuration: each tuple is 64-bit wide, 32-bit for each of the two fields, a pipeline throughput $P$ of 4 tuples per cycle, and a sorter capacity $k$ of 128 (supporting window sizes of up to $\textit{WS}=16K$ for the SWAG engine). The WS and WA values are configured at run time.
All presented designs are written in Verilog [^1] Available: (open source repository will be released after peer-review). The entities such as the compare-and-swap units for the sorter’s merger cells or the “$n$” entities of the prefix scan are parameterisable, such as with custom data widths and functionality, and are written by hand. The structures that use those entities including the reverse butterfly network and the parallel merge tree in the sorter are generated by scripts. These Verilog-generating scripts are written in Python to ease the inference of more complex or repetitive entity topologies for arbitrary $P$ values, as well as $k$ values for the sorter.
4.1. SWAG engine throughput and latency
[…]
Figure 10. Final input rate influenced by backpressure, $P=4$.
[…]
Figure 11. Simulated latency, $P=4$, $k=128$, $f_{\textit{target}}=125$ MHz
Figure 11 presents a similar experiment that measures the latency between the time the first tuple enters into the engine and the time there is a tuple in its output. As it can be observed, the WS does not influence the latency up to $\textit{WS}=128$, because the linear sorters in the sorter have a fixed latency ($k/P$) for an input to be extracted. For $\textit{WS}>k$ the latency scales linearly with WS, because the sorted windows are merged with a fixed rate $P$, and the merge phase is compulsory. WA does not impact the latency.
[…]
4.2. System-level evaluation
The resulting engines are evaluated on the Ultra96 development board with the Zynq UltraScale+ ZU3EG device, an MPSoC that runs Linux on its ARM A53 cores. This is an embedded platform, showcasing the resource-efficiency and adaptability of the proposed SWAG engine.
The Verilog code is implemented on the FPGA using AMD Vivado. The 2GB memory of the board is divided in two halves by the device tree, the kernel-mapped part, which is optimised for cached access to the CPU, and the non-kernel part, which is appropriate for FPGA use. The operating frequency of all designs is 250 MHz, with an internal frequency of 125 MHz, resulting in a 256-bit/cycle datapath. The critical path is coming from the sorter, which highlights the benefit of being able to alleviate a lower operating frequency of FPGAs with high-throughput processing. Still, 250 MHz is relatively close to saturating the memory bandwidth for this platform (Papaphilippou et al., 2020). Figure 12 visualises this experimental setup.
Figure 12. Experimental setup on the MPSoC platform.
Figure 13. SWAG engine speedup over one A53 ARM core for 3 operator types.
The data first pass through the high-throughput sorter, and its output is directed towards the proposed aggregation engines. The engine is either the group-by-aggregate version or the corresponding part of the SWAG engine.
The core uses a direct memory access (DMA) engine to transfer the data from the main memory to the sorter. The engine configuration such as the operator selection is managed by writing to a few memory-mapped registers. As soon as the data arrives to the sorter, they are immediately being sorted into chunks of up to $k$ size. The destination of those sorted chunks are either a temporary block memory (BRAM), the main memory, or the aggregation engine. Since the proposed solutions use a merge sorter, it is a recursive operation, and the engine only reads the sorter’s output on the last merging pass, hence the multiplexer.
The sorter does not currently need to output data to the main memory for the SWAG version, however, as the $k^{2}$-sized windows are deemed enough. Hence, the DRAM here is part of the evaluation platform and is not involved or needed in the calculation of the SWAG queries. The handling of the data by the sorter is preceded by a simple buffering mechanism that provides the studied windows in their entirety, as described in section 3.2.
The designs are compared against highly-optimised C++ implementations. The algorithms are relatively simple and they rely heavily on well-studied functions and structures from the C++ Standard Library. The GCC compilation flags are -O3 -march=native. These only target a single of the four available cores to avoid introducing more design variables. Due to Amdahl’s law and other inefficiencies coming from multi-processing, we can consider that the obtained FPGA speedups are at most roughly four times higher than multi-threaded implementations.
4.3. Group-by aggregation (preliminary experiment)
The proposed group-by-aggregate engine is adapted as a building block in the novel SWAG engine. This part of the evaluation preliminarily explores it as a standalone accelerator. The supported operations per $group$ ID in this implementation are: minimum, maximum, sum, count and distinct count.
The C++ baseline is equivalent to the functionality of the design, i.e. first sorting the data and then performing the aggregation serially. The sorting function is std::sort, which is a well regarded sorting baseline for generalisable use, although not the absolute fastest. The achieved speedup for $2^{16}$ tuples over the embedded processor varies between 22 to 28 according to the input distribution. This variation relates to the number of rows in the output. For instance, if all tuples have the same group ID, then the latency of interacting with the FPGA would be comparatively higher, as only a single tuple is produced as an output.
This speedup mostly comes from the sorter, as the studied CPU is already efficient for linearly skimming through a sorted stream while aggregating, also given its advantageous memory interfacing and speed. The speedup behaviour, such as for arbitrarily long streams, is inline with the sorter (Papaphilippou et al., 2020), hence not elaborating with comparison details. The advantages in resource utilisation are described in section 4.5, and the discussion of this standalone functionality in section 5.
4.4. Sliding window aggregation (SWAG)
The proposed SWAG engine is implemented as an AXI peripheral, and is used as a co-processor in our Linux-running evaluation environment. The supported operators are the ones of the latter plus the min./med./max. operator. It provides these aggregation results in the form of up to 3 tuples per group representing the minimum, median and maximum with the corresponding group ID. This is achieved by the proposed approach as introduced in section 3.2 that only leaves this subset to pass as valid after the first prefix scan in the pipeline. Since this is based on filtering to achieve aggregation, when there are any overlappings, such as with a singleton group whose minimum tuple is the maximum, such extracted tuples are considered common and the number of output tuples never surpasses the input tuples. In all other aggregation queries, the last tuple of each group is the one that gets updates to carry the aggregation summaries, as with the unmodified group-by aggregate engine of section 3.1.
The query selection is sum, distinct count and min./med./max. The logic behind this selection is to select a variation of algorithms that benefit the CPU and the FPGA at different degrees. For instance, the sum operator is incremental, and a pointer-based solution with software can potentially have an advantage over the FPGA solution, as it does not have to read entire windows for smaller window updates. On the other hand, min./med./max. is less trivial or rewarding to calculate incrementally on the CPU. The baseline software is based on C++ std::set and std::map to hold and update per-group information (notably similar performance to their hash-based unordered counterparts, not shown), and std::nth_element() to calculate the medians.
As demonstrated in figure 13, the FPGA advantage of the solution is different for the three operator types. There is a clear influence of the throughput behaviour, as was seen in the simulated SWAG engine of figure 10, since after $\textit{WS}>k=128$, there is the internal second pass to support the larger windows that halves the throughput. Still, for the majority of operators, WS and WA combinations, the three operators provide a speedup of one to two orders of magnitude. If we ignore the dip after the internal pass is used ($\textit{WS}<128$), we can see the upward trend in the figures with respect to the window size, which can be explained by the thoughtful use of the on-chip memory and longer memory transfers, as opposed to the CPU’s cache limitations. An exception is the combination of a high WS and low WA for the sum operator where the FPGA experiences a speedup lower than 1x. This comes from the software advantage of being able to efficiently do incremental updates for such a simple operator as sum, though according to a use case such as with 100% unique group IDs, WS=128, WA=64 the speedup is 164x (not shown).
The studied operators also behave differently under different input distributions at least when compared to the software equivalent. The input of sum and distinct count operators consists of 50% unique group IDs per group, which means every window has around WS/2 unique group IDs. The variations when changing from 100% unique to near 0% unique (single group ID) were relatively limited for these two operators, especially with regard to the FPGA performance, hence only the 50% version is shown to conserve space. For example, for WS=16384, WA=8192 the speedup on the distinct count operator is 317x, 420x and 476x for 100%, 50%, and single unique groups respectively. However, this is not the case for our main target min./med./max., hence the combined rightmost plot of figure 13. The output size varies according to the distribution, ranging from 1 to 3 tuples per group ID, according to the overlappings. If we select WS=16384, WA=64, the corresponding speedups for min./med./max. change to 59x, 239x and 472x. This is expected as the increased localities from having to operate on fewer elements can benefit software implementations when comparing the overheads of using a co-processor, while noting that the speedup is still impressive.
4.5. Resource utilisation
Scalability (synthesis-based)
The proposed aggregation engines are scalable to arbitrary degrees of parallelism $P$ to be able to saturate the bandwidth for current and future systems with a wide datapath. The total amount of data that can be produced or consumed by the engines per cycle is $P$ times data width. An exploration of the impact of $P$ and the data width on the resources is conducted with out-of-context synthesis results. The tuple (data) width is halved and used by the group ID and key fields for simplicity. The synthesis is done with Yosys (Wolf, 2016) for a more generic but modern 6LUT architecture, which can serve as high-level guidelines for the resource utilisation expectations in future applications.
Figure 14 summarises the results of this exploration for the group-by engine (upper) and the SWAG engine (lower), without the accompanying sorters and sliding window buffering. Since the first uses the unmodified open-source sorter, and the latter an adaptation of it, the readers can also refer to its original description (Papaphilippou et al., 2020) or an alternative implementation.
Figure 14. Synthesis results when varying tuple width and $P$ (sorter excluding).
For the ranges of interest, the registers and LUTs scale almost linearly with $P$ and the tuple width. The fastest growing component is the reverse butterfly network, with a complexity of $\Theta(P\log_{2}(P))$ switches (similar for the prefix scans), though the slight upward trend is not observed in these ranges yet. The register utilisation is always more for the SWAG engine than the group-by engine, and this can be explained by the decoupling of the simultaneous prefix scan part into two consecutive ones. This has lengthened the pipeline length, and thus, register use, as the presented implementations use a register per pipeline stage. Our use case has yielded a low enough register utilisation for our purposes, but future work can also focus on retiming to considerably reduce the register utilisation when necessary.
The LUT utilisation for the group-by engine and SWAG engine are similar, but there is variation between their corresponding configurations, as sometimes SWAG is more lightweight. Although the additional pipeline stages and the additional support of min./med./max. has an overhead in utilisation, many of the row/cardinality counters are narrower than the corresponding of the group-by engine. This is since the scope of the window size (WS) for the SWAG engine are limited to $k^{2}$ (here equal to $128^{2}$). The group-by engine uses counters of key-width, which also explains the higher impact from the tuple width in the upper plots of figure 14.
Overall, $P$ seems more influential to the LUT and register use than the tuple width, as it has the potential to scale to more orders of magnitude for the studied ranges. Even the highest points here ($P=64$, tuple width 256 bits) are feasible with today’s datacenter FPGAs, though any complications would first relate on how easy is to establish a sorted stream with such a high aggregate throughput.
Implementation
Table 1 introduces the look-up table (LUT) and flip-flop register (FF) utilisation as reported by Vivado 2024.1 (and 2020.1 for the unmodified sorter results for consistency with literature) for the selected target device. The presented utilisation is for the encapsulating AXI peripheral. Note that slight variations can occur when further exploring optimisation directives, interfacing, register placement etc., though they remain consistent between all presented runs.
Table 1. Vivado resource utilisation report summary
Design Group operators LUT FF
Sorter (Papaphilippou et al., 2020) N/A 34,101 47,122
Sorter & Group-by engine min., max., sum, count, distinct count 36,653 49,544
Adapted sorter1 & SWAG engine min./med./max. 41,963 54,437
Adapted sorter1 & SWAG engine min., max., sum, count, distinct count, min./med./max. 43,441 54,694
1with cardinality support, as introduced in section 3.3.
With respect to the group-by aggregate engine, we can see that there is about a 2.5K overhead on the number of Lookup-tables (LUTs) when becoming an add-on to an existing unmodified sorter. All engine designs here use a register per pipeline stage, which also introduced a 2.4K overhead on flip-flop (FF) registers for the engine.
The group-by engine essentially improves on the design of the PRRA-based distinct counter (Papaphilippou et al., 2023) of figure 1. If we compare its utilisation (2.55K LUT, 2.42K FF) to the unmodified PRRA (1.26K LUT, 2.29K FF (Papaphilippou et al., 2023)), it is easy to conclude that it uses at least 1.9x fewer FFs than any solution with two PRRAs, and similar LUTs before even adding the aggregation logic. This can be explained with the help of a simple model: when we abstract the node complexity to 1 for the butterfly network switches and all other entities, the complexity of the example PRRA-based engine of figure 1 decreases from $3P+2\times\mathrm{PRRA}$ to $12P+1\times\mathrm{PRRA}=2P\times\log_{2}(P)+P+$ 1 entities.
All changes to build the SWAG engine total around 7K LUTs and 5K FFs, and this includes the cardinality calculation support within the sorter for each tuple (see section 3.3). In terms of block RAM, it needs to accommodate $k^{2}$ tuples for the sliding window buffer and $k^{2}$ for the merge phase of the sorter, where $k$ is the sorter capacity. Since $k=128$, this totals $3k^{2}\times 64\textit{bits}=3\textit{Mibs}\approx 85\times 36\textit{Kibs}$, resulting in 89 BMEMs.
5. Related work
There is a moderate amount of research on database acceleration focusing on analytics (Papaphilippou and Luk, 2018). Similar to earlier works that accelerated joins (Casper and Olukotun, 2014; Owaida et al., 2017), there are tradeoffs when selecting between a sort and a hash-based solution (Moghaddamfar et al., 2023; Absalyamov et al., 2016), each with separate adaptation challenges (Mätas et al., 2022; Eryilmaz et al., 2021). For instance, even with a sorter-based pipeline, supporting arbitrary data in a join accelerator caused stalls and increased its memory requirements (Casper and Olukotun, 2014). This related to the existence of duplicates of unknown cardinality, which is somewhat reminiscent of median support, which affects the complexity of SWAG designs.
Group-by-aggregate operators have usually been a part of larger analytics designs (Lisa et al., 2018), as with the presented engine. This is due to the bandwidth limitations (Eryilmaz et al., 2021), which makes isolated solutions less practical. A group-by-aggregate-focused work demonstrated a speedup of 10x over the CPU (Absalyamov et al., 2016), but used multi-tasking, hashes and off-chip memory on a discontinued platform. A recent group-by-aggregator uses sorting to accelerate these operations (Xue et al., 2024), though it implements an SQL-based data management system, which is still inline with the observation on the scarcity of out-of-context group-by-aggregate accelerators.
On the other hand, SWAG engines exist stand-alone, also making direct comparisons more insightful. A state-of-the-art FPGA solution provides a notable speedup over prior work, but still operates with no ability to process multiple elements per cycle (Geethakumari and Sourdis, 2021b, 2023). It involves the implementation of a multi-level memory hierarchy based on off-chip memories (Geethakumari and Sourdis, 2021a), which is reflected on the complexity of the design. In contrast, our proposed SWAG engine is able to be evaluated on a resource-constrained embedded platform, also without the need for compression (Geethakumari and Sourdis, 2021b), while still having significant advantages.
Table 2. SWAG engine implementation comparison, inspired by (Geethakumari and Sourdis, 2023)
DFE(Ramakrishnan Geethakumari et al., 2019) MLQ(Geethakumari and Sourdis, 2021a) StreamZip(Geethakumari and Sourdis, 2021b, 2023) Proposed work
Group operators min./med./avg/max. min. min., max., sum, count, dist. count, min./med./max.
[…]
1Intel’s ALM is not directly comparable, but implies more logic per unit.
2All operators except summation also support IEEE-754 floating points.
[…]
6. Discussion and future work
The isolated task of group-by-aggregate (non-SWAG) is not necessarily expensive on other platforms, such as when using hash tables for smaller arrays instead of sorting on CPUs. Still, the engine simplifies the enablement of this functionality in programmable logic, such as for SWAG and high-throughput pipelined database analytics accelerators (Sun et al., 2021) to reduceunnecessary data transfers (Lisa et al., 2018).
The engines are highly customisable and parameterisable. The selection mechanism inside the SWAG engine can easily be generalised to support a percentile operator to select an arbitrary number of percentiles based on the pre-calculated cardinality. The average operation is supported in simulation, but is omitted for brevity, in order to avoid shifting the focus to floating point division optimisations. The throughput and latency limitations mentioned for SWAG in section 4.1 can be addressed. When a higher throughput is needed for large windows, a wider $P$ can be used, or two sorters can be combined and de-multiplexed using rate converters. A lower latency can be achieved by using lower $k$ values for $\textit{WS}\leq k$, or a higher $P$ for $\textit{WS}>k$. If the missing $\textit{WS}=256$ mode is needed, a build with $k=64$ or $256$ would suffice. Alternatively, this inherited limitation of the utilised sorter can be addressed at its source, to enable splitting the sorted sublists into smaller pieces to also conserve the merging throughput.
Other future work includes the automation of synthesising customised aggregation engines, such as by developing a high-level synthesis (HLS) library. Additional effort could be put into automating the inference of such circuits and modular components. The proposed aggregation engine could also be explored as a functional unit inside CPUs to work as a set of dedicated SIMD instructions.
7. Conclusions
The increasing memory datapath widths are becoming increasingly challenging to exploit with existing programming techniques and models. This includes naturally parallel problems such as aggregation. A modular and highly-adaptable engine is presented that uses prefix-scan and reverse butterfly topologies to perform a wide range of aggregation tasks with high-throughput as a non-blocking pipeline. It is then adapted to work as a highly competitive sliding window aggregation (SWAG) engine by incorporating a sorter that sorts each incoming buffered window. A wide selection of additions are proposed to an existing high-throughput sorter to append group cardinality information to the incoming tuples, in order to facilitate the calculation of the median operator. The resulting designs are shown to have a low hardware overhead, while supporting a superset of the functionality of alternatives. The SWAG engine exhibits up to hundreds of times of speedup when compared to a CPU-only approach on our embedded evaluation platform, and is able to process multiples of the input rate of the state-of-the-art on a datacenter grade FPGA. This is achieved without the need for off-chip memories.